OpenStudy (anonymous):

tan^2(60) + 4cos^2(45) - 9tan^2(30) - 8sin^2(30) The answer is zero but why!? i need step by step solution please

OpenStudy (anonymous):

why can't you just plug the whole thing into your calculator?

OpenStudy (anonymous):

because in the exam they already give you the answer, you have to do the procedure

OpenStudy (callisto):

tan60 = \(\sqrt{3}\) cos45 = \(\frac{1}{\sqrt2}\) tan30 = \(\frac{1}{\sqrt3}\) sin30 = 1/2 These are all special angles and you should remember them.

OpenStudy (anonymous):

yes i know them but i just continue to get 6 instead of 0

OpenStudy (anonymous):

I'm getting zero just fine though. note the squares.

OpenStudy (anonymous):

tan^2(60) = 3 ? 4cos^2(45) = 4 ?

OpenStudy (callisto):

4cos^2(45) = 4 (1/2) = ...?

OpenStudy (anonymous):

my math base is not good, that's the reason i wrote step by step solution

OpenStudy (callisto):

\[cos45 =\frac{1}{\sqrt2}\]\[cos^245 =(\frac{1}{\sqrt2})^2 = \frac{1}{2}\]\[4cos45 =4(\frac{1}{2}) = ...?\]

OpenStudy (anonymous):

4 (1/2) = 2

OpenStudy (callisto):

Yes.

OpenStudy (anonymous):

\[\frac{ \sin ^{2} (60)}{ \cos ^{2}(60)}+4\cos ^{2}(45)-9\frac{ \sin ^{2}(30) }{ \cos ^{2}(30) }-8\sin ^{2}(30)\]

OpenStudy (anonymous):

so, tan^2(60) = 3

OpenStudy (callisto):

tan^2(60) = 3 <- Yes

OpenStudy (anonymous):